A medical office does MRI (magnetic resonance imagingl scans on patients. it has two MRI mechines. Each scan requires a technician to sun the MRI machine Below is the production function for the offic which shows how many patients the techniclans can scan 10 14 Supoose that the price of the lobor input the MRi technician) ls $200 per day in wages and benelts Suppose also thet the capital input Ohe MRI mechinesj hes a price of $150 per day for each machine, on long term lense ) To do 10 scans in a dey, how mary technicians are needed?2echnicians b) What is the lsbor cost of doing 10 scans in a day? What is the total cost of doing 10 scens per day (ossuning there ae two Mie machines in the office Now suppose that the number of scans inceases to 1 di What is the variable cost of doing 17 scens in a day? S 2000 2.30o el What is the fixed cost of doing 17 scans in a day?
Answer #1

Wages = $200 per day

Rent= $150 per day

(a) To do 10 scans in a day, 2 technicians are needed.

(b) Labor cost of doing 10 scans in a day = wL = (200)(2)= $400

(c) Total cost of doing 10 scans per day (two MRI machines in the office)= wL + rK = (200)(2)+ (150)(2)= 400+300= $700

(d) Number of scans increases to 17. To do 17 scans in a day , 4 technicians are needed.

Variable cost =wL= (200)(4)= $800

(e) Fixed cost = (150)(2) = $300.